# Distributed High-Dimensional Quantile Regression: Estimation Efficiency and Support Recovery

### v1

@@ 4 Simulation Studies
In this section, we provide simulation studies to assess the performance of our DHSQR estimator. We generate synthetic data from the following linear models, corresponding to the homoscedastic error case (model 1) and the heteroscedastic error case (model 2):
@@
For the choice of the kernel function, we use the standard Gaussian kernel function that satisfies Assumption 3.1. For the global and local bandwidth, we set $h=5(s\log N/n)^{1/3}$ and $b=0.53(s\log n/n)^{1/3}$ , respectively, according to the theoretical results in Theorem 3.7 and 3.8. The regularization parameters $\lambda_{N,g}$ are selected by cross-validation. Specifically, we choose $C_{0}$ to minimize the check loss on the validation set. All the simulation results are the average of 100 independent experiments.
@@
which is commonly used as an evaluation of support recovery. Note that $F_{1}$-score$=1$ implies perfect support recovery.
We compare the finite-sample performance of the DHSQR estimator and the other four estimators:
@@ 4.1 Effect of the number of iterations Under Heavy-Tailed Noise
We first show the effect of the number of iterations in our proposed method. We fix the total sample size $N=20000$ and local sample size $n=500$. We plot the $\ell_{2}$-error from the true QR coefficients versus the number of iterations. Since the Avg-DC only requires one-shot communication, we use a horizontal line to show its performance. Since the Avg-DC only requires one-shot communication, we use a horizontal line to show its performance. The results are shown in Figure 1.
Figure 1: The $\ell_{2}$-error with an error bound between the true parameter and the estimated parameter versus the number of iterations($\tau=0.5$). In the left panel, from top to bottom represent noise distributions that are Normal, $t_{3}$, and Cauchy distribution for the homoscedastic error case, respectively. In the right panel, from top to bottom represent noise distributions as Normal, $t_{3}$, and Cauchy distribution for the heteroscedastic error case, respectively.
From the result, our DHSQR estimator outperforms the Avg-DC and DPQR estimators in all the cases after a few iterations, and the estimated $\ell_{2}$-error is very close to that of the pooled DHSQR estimator. Our method is robust for different noise settings, while DPQR performs poorly for Heavy-Tailed Noise. Interestingly, for the heteroscedastic error case, the DREL estimator becomes unstable and fails to converge since it ignores the dependence between the error term and covariate. These results confirm our theoretical findings in section 3. For the rest of the experiments in this section, we fix the number of iterations $T=10$.
@@ 4.2 Effect of Total Sample Size and Local Sample Size
We investigate the performance of our proposed estimator under varying total and local sample sizes. In this section, we only consider the setting with normal error case and quantile level $\tau=0.5$. For the effect of total sample size, we the local sample size $n=500$, and vary the total sample size $N\in\{5000,10000,15000,20000,25000\}$. For the effect of local sample size, we fix the total sample size $N=20000$, and vary the local sample size $n\in\{200,500,1000,2000\}$. We plot the $\ell_{2}$-error and $F_{1}$-score of the five estimators under different models in Figure 2 and 3.
The $\ell_{2}$-error of all methods decreases with an increase in the number of total sample size $N$ and local sample size $n$, as shown in Figure 2. For DHSQR, DPQR, and DREL estimators, as $N$ increases or $n$ increases, the $\ell_{2}$-error tends to decrease, and both of them outperform the Avg-DC estimator. We can see that the $\ell_{2}$-error of DHSQR is very close to that of the Pooled DHSQR estimator in both homoscedastic and heteroscedastic error cases. The DHSQR estimator is better than the other three distributed estimator. In the heteroscedastic error case, the DREL estimator’s $\ell_{2}$-error is significantly worse compared to DHSQR and DPQR. The Pooled DHSQR estimator’s performance is not significantly affected by variations in $n$, while Avg-DC shows minimal sensitivity to the changes in $N$.
@@
Figure 3: The $F_{1}$ score from the true parameter versus the number of total and local sample size with a fixed quantile level $\tau=0.5$. (a) and (b) shows the effect of different total sample sizes $N$ for the homoscedastic and heteroscedastic error cases, respectively. (c) and (d) shows the effect of different local sample sizes $n$ for the homoscedastic and heteroscedastic error cases, respectively.
Due to the space limit, we have also included additional real data application and simulated experiments for sensitivity analysis for the bandwidth, computation time and different quantile levels in Appendix C.
@@ 5 Conclusion
In this paper, we propose an efficient distributed quantile regression method for dealing with heterogeneous data. By constructing a kernel-based pseudo covariate and response, we transform the non-smooth quantile regression problem into a smooth least square problem. Based on this procedure, we establish a double-smoothing surrogate likelihood framework to facilitate distributed learning. An efficient algorithm is developed, which enjoys computation and communication efficiency. Both theoretical analysis and simulation studies demonstrated the superior performance of the proposed algorithm. As future work, we will investigate the inference problem in distributed high-dimensional quantile regression based on the debiased procedure (Zhao et al., 2014; Bradic & Kolar, 2017). It is also interesting to develop a decentralized distributed algorithm for quantile regression over a network data structure.
@@ Acknowledgments
The authors thank the area chair and three anonymous referees for their constructive suggestions, which significantly improve this article.
Impact Statements
@@ A.1 Some Technical Lemmas
Let $\xi_{1},\ldots,\xi_{N}$ be independent random variables with zero mean. Suppose that there exists some $t>0$ and $\bar{B}_{N}$ such that $\sum_{i=1}^{N}E\xi_{i}^{2}\exp\left(t\left|\xi_{i}\right|\right)\leq\bar{B}_{N}^{2}$. Then for any $0<x\leq\bar{B}_{N}$,
@@
$\underset{j\leq 5^{s}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|\leq\underset{j\leq 5^{s}}{\operatorname{sup}}\underset{\|\boldsymbol{\beta}_{S}-\boldsymbol{\beta}^{*}_{S}\|\leq a_{N}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}(\boldsymbol{\beta})-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|.$ (23)
@@
$\underset{j\leq 5^{s}}{\operatorname{sup}}\underset{\|\boldsymbol{\beta}_{S}-\boldsymbol{\beta}^{*}_{S}\|\leq a_{N}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}(\boldsymbol{\beta})-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|-\underset{j\leq 5^{s}}{\operatorname{sup}}\underset{1\leq k\leq N^{C_{1}s}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}(\boldsymbol{\beta}_{k})-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|$
@@
$\underset{j\leq 5^{s}}{\operatorname{sup}}\underset{\|\boldsymbol{\beta}_{S}-\boldsymbol{\beta}^{*}_{S}\|\leq a_{N}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}(\boldsymbol{\beta})-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|-\underset{j\leq 5^{s}}{\operatorname{sup}}\underset{1\leq k\leq N^{c_{1}s}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}(\boldsymbol{\beta}_{k})-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|=\mathcal{O}_{\mathbb{P}}(N^{-\gamma})$ (24)
@@
$=\boldsymbol{X}_{i,S}\boldsymbol{X}_{i,S}^{\mathrm{T}}\left(f_{\varepsilon\mid\boldsymbol{X}}(0)+O\left(h+\boldsymbol{X}_{i,S}^{\mathrm{T}}\left(\boldsymbol{\beta}_{S}-\boldsymbol{\beta}_{S}^{*}\right)\right)\right).$
@@
$\underset{1\leq k\leq N^{C_{1}s}}{\operatorname{sup}}{\mathbb{P}}\left(\left|\sum_{i=1}^{N}\left(\xi_{i,j,k}-\mathbb{E}\xi_{i,j,k}\right)\right|\geq C\sqrt{Nhs\log N}\right)=O\left(N^{-\gamma s}\right).$
@@
$\max_{1\leq j\leq p}\left\|\frac{1}{N}\sum_{i=1}^{N}\left|X_{ij}\right|\boldsymbol{X}_{i,S}\boldsymbol{X}_{i,S}^{\mathrm{T}}\right\|_{\mathrm{op}}=\mathcal{O}_{\mathbb{P}}(1)$
@@
$\max_{1\leq j\leq p}\left\|\frac{1}{N}\sum_{i=1}^{N}\left|X_{ij}\right|\boldsymbol{X}_{i,S}\boldsymbol{X}_{i,S}^{\mathrm{T}}\right\|_{\mathrm{op}}\leq\max_{1\leq j\leq p}\max_{1\leq k\leq 5^{s}}\frac{1}{N}\sum_{i=1}^{N}\left|X_{ij}\right|\left(\boldsymbol{v}_{k}^{\mathrm{T}}\boldsymbol{X}_{i,S}\right).$
Denote $\widetilde{X}_{ij}=X_{ij}\mathbf{1}\left(\left|X_{ij}\right|\leq\log N\right)$, under Assumption 3.3, we can alternatively show that
@@
Set $Y_{ijk}=\left|\widetilde{X}_{ij}\right|\left(\boldsymbol{v}_{k}^{\mathrm{T}}\boldsymbol{X}_{i,\mathcal{S}}\right)^{2}\mathbf{1}\left\{\left|\widetilde{X}_{ij}\right|\left(\boldsymbol{v}_{k}^{\mathrm{T}}\boldsymbol{X}_{i,\mathcal{S}}\right)^{2}\geq(s+1)\log^{3}N\right\}$, note that
$Np5^{s}\max_{i,j,k}{\mathbb{P}}\left(\left|\widetilde{X}_{ij}\right|\left(\boldsymbol{v}_{k}^{\mathrm{T}}\boldsymbol{X}_{i,\mathcal{S}}\right)^{2}\mathbf{1}\left\{\left|\widetilde{X}_{ij}\right|\left(\boldsymbol{v}_{k}^{\mathrm{T}}\boldsymbol{X}_{i,\mathcal{S}}\right)^{2}\geq(s+1)\log^{3}N\right\}\right)=\mathcal{O}(1),$
@@
$Np9^{s}\left(e^{-q_{1}N}+e^{-q_{2}\frac{N}{(s+1)\log^{3}n}}\right)=\mathcal{O}(1).$
@@
A.2 Proof of Proposition 1 and Propositin 2
We first give some propositions to prove Theorem 3.7 and Theorem 3.8.
@@ A.2 Proof of Proposition 1 and Propositin 2
For the convex quadratic optimization (11) in the main text, define $\mathcal{A}=\left\{\delta\in\mathbb{R}^{p}:|\delta|_{1}\leq 4s^{1/2}|\delta|_{2}\right\}$, if the following conditions
@@
$\left|f_{\varepsilon|\boldsymbol{X}}(0)\widehat{\boldsymbol{\Sigma}}-\boldsymbol{I}\right|_{\infty}\leq C\sqrt{\frac{\log p}{N}}$
@@
$\left|\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{X}_{i}\left\{F_{\varepsilon|\boldsymbol{X}}(\boldsymbol{X}_{i,S}^{\mathrm{T}}(\widehat{\boldsymbol{\beta}}_{0,S}-\boldsymbol{\beta}_{S}^{*}))-F_{\varepsilon|\boldsymbol{X}}(0)\right\}-\boldsymbol{I}\left(\widehat{\boldsymbol{\beta}}_{0}-\boldsymbol{\beta}^{*}\right)\right|_{\infty}=\mathcal{O}_{\mathbb{P}}\left(\sqrt{\frac{s\log p}{N}}a_{N}+a_{N}^{2}\right)$ (28)
@@
$\left|(\widehat{\boldsymbol{D}}_{h}-\boldsymbol{I})(\widehat{\boldsymbol{\beta}}_{0}-\boldsymbol{\beta}^{*})\right|_{\infty}$ $\leq\left\|(\widehat{\boldsymbol{D}}_{h,S\times S}-\boldsymbol{I}_{S\times S})\right\|_{op}\left|(\widehat{\boldsymbol{\beta}}_{0,S}-\boldsymbol{\beta}^{*}_{S})\right|_{2}=\mathcal{O}_{\mathbb{P}}\left(a_{N}\sqrt{\frac{s\log N}{Nh}}+a_{N}^{2}+a_{N}h\right)$ (30)
@@
This complete the proof. ∎
@@
$\left|\boldsymbol{A}\boldsymbol{\beta}^{*}-\boldsymbol{a}\right|_{\infty}$ $=\left|\widehat{\boldsymbol{D}}_{1,b}\boldsymbol{\beta}^{*}-\boldsymbol{z}_{N}-\left(\widehat{\boldsymbol{D}}_{1,b}-\widehat{\boldsymbol{D}}_{h}\right)\widehat{\boldsymbol{\beta}}_{0}\right|_{\infty}\leq\left|\widehat{\boldsymbol{D}}_{h}\boldsymbol{\beta}^{*}-\boldsymbol{z}_{N}\right|_{\infty}+\left|\left(\widehat{\boldsymbol{D}}_{1,b}-\widehat{\boldsymbol{D}}_{h}\right)\left(\widehat{\boldsymbol{\beta}}_{0}-\boldsymbol{\beta}^{*}\right)\right|_{\infty}$ (32)
@@
We use the primal dual witness (PWD) construction from (Wainwright, 2009) to prove Theorem 3.9. By Karush-Kuhn-Tucher’s Theorem, there exists subgradient $\widehat{\boldsymbol{Z}}=\partial|\widehat{\boldsymbol{beta}}_{1}|$ such that
@@
    $\widehat{\boldsymbol{D}}_{1,b,S\times S}\widetilde{\boldsymbol{\beta}}_{S}-\left\{z_{N}+\left(\widehat{\boldsymbol{D}}_{1,b}-\widehat{\boldsymbol{D}}_{h}\right)\widehat{\boldsymbol{\beta}}_{0}\right\}_{S}+\lambda_{N}\widetilde{\boldsymbol{Z}}_{S}=0$ (35)
@@
$\boldsymbol{M}_{2}=\widehat{\boldsymbol{D}}_{1,b,S^{c}\times S}\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}\widetilde{Z}_{S},\quad\boldsymbol{M}_{3}=\lambda_{N}^{-1}\left\{\boldsymbol{z}_{N}-\widehat{\boldsymbol{D}}_{h}\beta^{*}\right\}_{S^{c}},$
@@
By (34) in the proof of Theorem 3.7, let $C_{0}$ be sufficient large, we have
$\left|-\lambda_{N}^{-1}\left(\left\{\boldsymbol{z}_{N}-\widehat{\boldsymbol{D}}_{h}\boldsymbol{\beta}^{*}\right\}_{S}+\left(\widehat{\boldsymbol{D}}_{h,S\times\{1,\ldots,p+1\}}-\widehat{\boldsymbol{D}}_{1,b,S\times\{1,\ldots,p+1\}}\right)\left(\boldsymbol{\beta}^{*}-\widehat{\boldsymbol{\beta}}_{0}\right)\right)\right|_{\infty}\leq\alpha/12$ (37)
@@
$\left\|\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}-\boldsymbol{I}_{S\times S}^{-1}\right\|_{op}\leq\left\|\boldsymbol{I}_{S\times S}^{-1}\right\|_{op}^{2}\left\|\widehat{\boldsymbol{D}}_{1,b,S\times S}-\boldsymbol{I}_{S\times S}\right\|_{op},$

### v3

@@ 4 Simulation Studies
In this section, we provide simulation studies to assess the performance of our DHSQR estimator  [^1] The R code to reproduce our experimental results is available in https://github.com/WangCaixing-96/DHSQR.. We generate synthetic data from the following linear models, corresponding to the homoscedastic error case (model 1) and the heteroscedastic error case (model 2):
@@
For the choice of the kernel function, we use the standard Gaussian kernel function that satisfies Assumption 3.1. For the global and local bandwidth, we set $h=5(s\log N/n)^{1/3}$ and $b=0.53(s\log n/n)^{1/3}$ , respectively, according to the theoretical results in Theorem 3.7 and 3.8. The regularization parameters $\lambda_{N,g}$ are selected by validation. Specifically, we choose $C_{0}$ to minimize the check loss on the validation set. All the simulation results are the average of 100 independent experiments.
Figure 1: The $\ell_{2}$-error with an error bound between the true parameter and the estimated parameter versus the number of iterations with a fixed quantile level $\tau=0.5$. In the left panel, from top to bottom represent noise distributions that are Normal, $t_{3}$, and Cauchy distribution for the homoscedastic error case, respectively. In the right panel, from top to bottom represent noise distributions as Normal, $t_{3}$, and Cauchy distribution for the heteroscedastic error case, respectively.
@@
which is commonly used as an evaluation of support recovery. Note that $F_{1}$-score$=1$ implies perfect support recovery. We compare the finite sample performance of the DHSQR estimator and the other four estimators:
@@ 4.1 Effect of the number of iterations Under Heavy-Tailed Noise
We first show the effect of the number of iterations in our proposed method. We fix the total sample size $N=20000$ and local sample size $n=500$. We plot the $\ell_{2}$-error from the true QR coefficients versus the number of iterations. Since the Avg-DC only requires one-shot communication, we use a horizontal line to show its performance. The results are shown in Figure 1.
From the result, our DHSQR estimator outperforms the Avg-DC and DPQR estimators in all the cases after a few iterations, and the estimated $\ell_{2}$-error is very close to that of the pooled DHSQR estimator. Our method is robust for different noise settings, while DPQR performs poorly for heavy-tailed noise. Interestingly, for the heteroscedastic error case, the DREL estimator becomes unstable and fails to converge since it ignores the dependence between the error term and covariates. These results confirm our theoretical findings in section 3. For the rest of the experiments in this section, we fix the number of iterations $T=10$.
@@ 4.2 Effect of Total Sample Size and Local Sample Size
In this section, we investigate the performance of our proposed estimator under varying total and local sample sizes. We only consider the setting with normal error case and quantile level $\tau=0.5$. For the effect of total sample size, we fix the local sample size $n=500$, and vary the total sample size $N\in\{5000,10000,15000,20000,25000\}$. For the effect of local sample size, we fix the total sample size $N=20000$, and vary the local sample size $n\in\{200,500,1000,2000\}$. We plot the $\ell_{2}$-error and $F_{1}$-score of the five estimators under different models in Figure 2 and 3.
Figure 3: The $F_{1}$ score from the true parameter versus the number of total and local sample size with a fixed quantile level $\tau=0.5$. In the top panel, from left to right show the effect of different total sample sizes $N$ for the homoscedastic and heteroscedastic error cases, respectively. In the bottom panel, from left to right show the effect of different local sample sizes $n$ for the homoscedastic and heteroscedastic error cases, respectively.
The $\ell_{2}$-error of all methods decreases with an increase in the total sample size $N$ and local sample size $n$, as shown in Figure 2. For DHSQR, DPQR, and DREL estimators, as $N$ increases or $n$ increases, the $\ell_{2}$-error tends to decrease, and both of them outperform the Avg-DC estimator. We can see that the $\ell_{2}$-error of DHSQR is very close to that of the Pooled DHSQR estimator in both homoscedastic and heteroscedastic error cases. The DHSQR estimator is better than other three distributed estimators. In the heteroscedastic error case, the DREL estimator’s $\ell_{2}$-error is significantly worse compared to DHSQR and DPQR. The Pooled DHSQR estimator’s performance is not significantly affected by variations in $n$, while Avg-DC shows minimal sensitivity to the changes in $N$.
@@
Due to the space limit, we have also included an additional real data application and simulated experiments for sensitivity analysis for the bandwidth, computation time, different quantile levels, and different settings of parameters $\boldsymbol{\beta}^{*}$ in Appendix C.
@@ 5 Conclusion
In this paper, we propose an efficient distributed quantile regression method for dealing with heterogeneous data. By constructing a kernel-based pseudo covariate and response, we transform the non-smooth quantile regression problem into a smooth least squares problem. Based on this procedure, we establish a double-smoothing surrogate likelihood framework to facilitate distributed learning. An efficient algorithm is developed, which enjoys computation and communication efficiency. Both theoretical analysis and simulation studies demonstrated the superior performance of the proposed algorithm. As future work, we will investigate the inference problem in distributed high-dimensional quantile regression based on the debiased procedure (Zhao et al., 2014; Bradic & Kolar, 2017). It is also interesting to develop a decentralized distributed algorithm for quantile regression over a network data structure.
@@ Acknowledgments
The authors thank the area chair and the anonymous referees for their constructive suggestions, which significantly improved this article. This research is supported in part by NSFC-12371270 and Shanghai Science and Technology Development Funds (23JC1402100). This research is also supported by Shanghai Research Center for Data Science and Decision Technology.
Impact Statement
@@ A.1 Some Technical Lemmas
Let $\xi_{1},\ldots,\xi_{N}$ be independent random variables with zero mean. Suppose that there exists some $t>0$ and $\bar{B}_{N}$ such that $\sum_{i=1}^{N}\mathbb{E}\xi_{i}^{2}\exp\left(t\left|\xi_{i}\right|\right)\leq\bar{B}_{N}^{2}$. Then for any $0<x\leq\bar{B}_{N}$,
@@
$\underset{j\leq 5^{s}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|\leq\underset{j\leq 5^{s}}{\operatorname{sup}}\underset{|\boldsymbol{\beta}_{S}-\boldsymbol{\beta}^{*}_{S}|_{2}\leq a_{N}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}(\boldsymbol{\beta})-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|.$ (23)
@@
$\underset{j\leq 5^{s}}{\operatorname{sup}}\underset{|\boldsymbol{\beta}_{S}-\boldsymbol{\beta}^{*}_{S}|_{2}\leq a_{N}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}(\boldsymbol{\beta})-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|-\underset{j\leq 5^{s}}{\operatorname{sup}}\underset{1\leq k\leq N^{C_{1}s}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}(\boldsymbol{\beta}_{k})-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|$
@@
$\underset{j\leq 5^{s}}{\operatorname{sup}}\underset{|\boldsymbol{\beta}_{S}-\boldsymbol{\beta}^{*}_{S}|_{2}\leq a_{N}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}(\boldsymbol{\beta})-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|-\underset{j\leq 5^{s}}{\operatorname{sup}}\underset{1\leq k\leq N^{c_{1}s}}{\operatorname{sup}}\left|\boldsymbol{v}_{j}^{\mathrm{T}}\{\widehat{\boldsymbol{D}}_{h,S\times S}(\boldsymbol{\beta}_{k})-\boldsymbol{I}_{S\times S}\}\boldsymbol{v}_{j}\right|=\mathcal{O}_{\mathbb{P}}(N^{-\gamma}).$ (24)
@@
$=\boldsymbol{X}_{i,S}\boldsymbol{X}_{i,S}^{\mathrm{T}}\left(f_{\varepsilon\mid\boldsymbol{X}}(0)+\mathcal{O}\left(h+\boldsymbol{X}_{i,S}^{\mathrm{T}}\left(\boldsymbol{\beta}_{S}-\boldsymbol{\beta}_{S}^{*}\right)\right)\right).$
@@
$\underset{1\leq k\leq N^{C_{1}s}}{\operatorname{sup}}{\mathbb{P}}\left(\left|\sum_{i=1}^{N}\left(\xi_{i,j,k}-\mathbb{E}\xi_{i,j,k}\right)\right|\geq C\sqrt{Nhs\log N}\right)=\mathcal{O}\left(N^{-\gamma s}\right).$
@@
$\max_{1\leq j\leq p}\left\|\frac{1}{N}\sum_{i=1}^{N}\left|X_{ij}\right|\boldsymbol{X}_{i,S}\boldsymbol{X}_{i,S}^{\mathrm{T}}\right\|_{op}=\mathcal{O}_{\mathbb{P}}(1).$
@@
$\max_{1\leq j\leq p}\left\|\frac{1}{N}\sum_{i=1}^{N}\left|X_{ij}\right|\boldsymbol{X}_{i,S}\boldsymbol{X}_{i,S}^{\mathrm{T}}\right\|_{op}\leq\max_{1\leq j\leq p}\max_{1\leq k\leq 5^{s}}\frac{1}{N}\sum_{i=1}^{N}\left|X_{ij}\right|\left(\boldsymbol{v}_{k}^{\mathrm{T}}\boldsymbol{X}_{i,S}\right).$
Denote $\widetilde{X}_{ij}=X_{ij}I\left(\left|X_{ij}\right|\leq\log N\right)$, under Assumption 3.3, we can alternatively show that
@@
Set $Y_{ijk}=\left|\widetilde{X}_{ij}\right|\left(\boldsymbol{v}_{k}^{\mathrm{T}}\boldsymbol{X}_{i,\mathcal{S}}\right)^{2}I\left\{\left|\widetilde{X}_{ij}\right|\left(\boldsymbol{v}_{k}^{\mathrm{T}}\boldsymbol{X}_{i,\mathcal{S}}\right)^{2}\geq(s+1)\log^{3}N\right\}$, note that
$Np5^{s}\max_{i,j,k}{\mathbb{P}}\left(\left|\widetilde{X}_{ij}\right|\left(\boldsymbol{v}_{k}^{\mathrm{T}}\boldsymbol{X}_{i,\mathcal{S}}\right)^{2}I\left\{\left|\widetilde{X}_{ij}\right|\left(\boldsymbol{v}_{k}^{\mathrm{T}}\boldsymbol{X}_{i,\mathcal{S}}\right)^{2}\geq(s+1)\log^{3}N\right\}\right)=o(1),$
@@
$Np9^{s}\left(e^{-q_{1}N}+e^{-q_{2}\frac{N}{(s+1)\log^{3}n}}\right)=o(1).$
@@
A.2 Proof of Proposition A.4 and Proposition A.5
We first give some propositions to prove Theorem 3.7 and Theorem 3.8. Consider a general convex quadratic optimization as follows,
$\widehat{\boldsymbol{\beta}}=\underset{\boldsymbol{\beta}\in\mathbb{R}^{p+1}}{\mathop{\rm arg}\min}\frac{1}{2}\boldsymbol{\beta}^{\mathrm{T}}\boldsymbol{A}\boldsymbol{\beta}-\boldsymbol{\beta}^{\mathrm{T}}\boldsymbol{a}+\lambda_{N}|\boldsymbol{\beta}|_{1},$ (27)
where $\boldsymbol{A}$ is a non-negative definite matrix and $\boldsymbol{a}$ is a vector in $\mathbb{R}^{p+1}$, then we can prove the following proposition.
@@
For the convex quadratic optimization in (27), define $\mathcal{A}=\left\{\delta\in\mathbb{R}^{p}:|\delta|_{1}\leq 4s^{1/2}|\delta|_{2}\right\}$, if the following conditions
@@
$\left|f_{\varepsilon|\boldsymbol{X}}(0)\widehat{\boldsymbol{\Sigma}}-\boldsymbol{I}\right|_{\infty}\leq C\sqrt{\frac{\log p}{N}}.$
@@
$\left|\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{X}_{i}\left\{F_{\varepsilon|\boldsymbol{X}}(\boldsymbol{X}_{i,S}^{\mathrm{T}}(\widehat{\boldsymbol{\beta}}_{0,S}-\boldsymbol{\beta}_{S}^{*}))-F_{\varepsilon|\boldsymbol{X}}(0)\right\}-\boldsymbol{I}\left(\widehat{\boldsymbol{\beta}}_{0}-\boldsymbol{\beta}^{*}\right)\right|_{\infty}=\mathcal{O}_{\mathbb{P}}\left(\sqrt{\frac{s\log p}{N}}a_{N}+a_{N}^{2}\right).$ (29)
@@
$\left|(\widehat{\boldsymbol{D}}_{h}-\boldsymbol{I})(\widehat{\boldsymbol{\beta}}_{0}-\boldsymbol{\beta}^{*})\right|_{\infty}$ $\leq\left\|(\widehat{\boldsymbol{D}}_{h,S\times S}-\boldsymbol{I}_{S\times S})\right\|_{op}\left|(\widehat{\boldsymbol{\beta}}_{0,S}-\boldsymbol{\beta}^{*}_{S})\right|_{2}=\mathcal{O}_{\mathbb{P}}\left(a_{N}\sqrt{\frac{s\log N}{Nh}}+a_{N}^{2}+a_{N}h\right).$ (31)
@@
This completes the proof. ∎
@@
$\left|\boldsymbol{A}\boldsymbol{\beta}^{*}-\boldsymbol{a}\right|_{\infty}$ $=\left|\widehat{\boldsymbol{D}}_{1,b}\boldsymbol{\beta}^{*}-\boldsymbol{z}_{N}-\left(\widehat{\boldsymbol{D}}_{1,b}-\widehat{\boldsymbol{D}}_{h}\right)\widehat{\boldsymbol{\beta}}_{0}\right|_{\infty}\leq\left|\widehat{\boldsymbol{D}}_{h}\boldsymbol{\beta}^{*}-\boldsymbol{z}_{N}\right|_{\infty}+\left|\left(\widehat{\boldsymbol{D}}_{1,b}-\widehat{\boldsymbol{D}}_{h}\right)\left(\widehat{\boldsymbol{\beta}}_{0}-\boldsymbol{\beta}^{*}\right)\right|_{\infty}.$ (33)
@@
We use the primal dual witness (PWD) construction from (Wainwright, 2009) to prove Theorem 3.9. By Karush-Kuhn-Tucker’s Theorem, there exists subgradient $\widehat{\boldsymbol{Z}}$ such that
@@
    $\widehat{\boldsymbol{D}}_{1,b,S\times S}\widetilde{\boldsymbol{\beta}}_{S}-\left\{z_{N}+\left(\widehat{\boldsymbol{D}}_{1,b}-\widehat{\boldsymbol{D}}_{h}\right)\widehat{\boldsymbol{\beta}}_{0}\right\}_{S}+\lambda_{N}\widetilde{\boldsymbol{Z}}_{S}=0.$ (36)
@@
$\boldsymbol{M}_{2}=\widehat{\boldsymbol{D}}_{1,b,S^{c}\times S}\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}\widetilde{\boldsymbol{Z}}_{S},\quad\boldsymbol{M}_{3}=\lambda_{N}^{-1}\left\{\boldsymbol{z}_{N}-\widehat{\boldsymbol{D}}_{h}\boldsymbol{\beta}^{*}\right\}_{S^{c}},$
@@
By (35) in the proof of Theorem 3.7, let $C_{0}$ be sufficiently large, we have
$\left|-\lambda_{N}^{-1}\left(\left\{\boldsymbol{z}_{N}-\widehat{\boldsymbol{D}}_{h}\boldsymbol{\beta}^{*}\right\}_{S}+\left(\widehat{\boldsymbol{D}}_{h,S\times\{1,\ldots,p+1\}}-\widehat{\boldsymbol{D}}_{1,b,S\times\{1,\ldots,p+1\}}\right)\left(\boldsymbol{\beta}^{*}-\widehat{\boldsymbol{\beta}}_{0}\right)\right)\right|_{\infty}\leq\alpha/12,$ (38)
@@
$\left\|\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}-\boldsymbol{I}_{S\times S}^{-1}\right\|_{op}\leq\left\|\boldsymbol{I}_{S\times S}^{-1}\right\|_{op}^{2}\left\|\widehat{\boldsymbol{D}}_{1,b,S\times S}-\boldsymbol{I}_{S\times S}\right\|_{op}.$